OF M A F12KN نررر منر ه يع تی 7CP VIEW

The beam is 132kg/m

What is the average shear stress on each pin?

Total mass = 132kg/m * 12m = 1584 kg

P = 15,539 N explain how

Reaction 1 = 7770 N + 300 N , Va = Ra = 8070 N

Reaction 2 = 7770 N + 900 N , Vb = Rb = 8670 N

Show the steps for calculating the static reactions.

____________________________________________________________________

What is the smallest diameter A36 steel pin that can be used?

A36yield = 250 MPa -> \sigma = \frac{F}{A}   F A 8670N 250 106.00035m

тD А Тр? А. 4   .000035 4 A* 4 D = .0067m 1000 6.7mm TT T  

What is the smallest diameter 2014 Aluminum pin that can be used with a 1.5 Factor of Safety (F.S.)?

8670 N * 1.5 = 13005 N   F 13005N .0000756m A 172 106

0000756 4 A* 4 D = .0098m 1000 9.8mm TT T
Answer #1

45 600 RB 1.2 KN (19210) L Lis8xg) Torul ut= 15840 N <15.8YUN HA 1.2cosuso RBlos 60 P T Rosinbo VA 1.2 sinuso {=0 H = 1.2 coTalling moment about C +3.96x1.5 - Fashy cox 30 VAX 9 +41-88x4.5), VA= 47.52 + 2.55RB VA= 0.28883 + 5.28 using this in (1) 0.